OpenStudy (anonymous):

Find the diameter of a sphere with volume 927pi inches to the 3rd

OpenStudy (anonymous):

do you know the formula for the volume of a circle??

OpenStudy (anonymous):

well because you were looking for the formula online i will teach you the way to work this problem out one by one. So because you have the formula what do you think you have to do first?

OpenStudy (anonymous):

I have the answer

OpenStudy (anonymous):

I fyou don't know you will need to write down the formula first then the volume of the circle like this \[4/3\pi(r)^3=927\pi\]

OpenStudy (anonymous):

then you will need to take out the fraction and the \[\pi\] so you will end up here if you did it correctly \[r^3 = 1236\]

OpenStudy (anonymous):

then the radius will be 5.929.......

OpenStudy (anonymous):

then multiply the radius by 2 and you will get 11.8.............

OpenStudy (anonymous):

11.8 will be your diameter